I need to know the answer and how you found the answer of 5x-3=4 and 2x+8=14-x
Answer by stanbon(75887)

I need to know the answer and how you found the answer of
5x-3=4
Add 3 to both sides to get:
5x = 7
Divide both sides by 5 to get:
x = 7/5 or 1.4
---------------------------

2x+8=14-x
Add x to both sides to get:
3x + 8 = 14
Subtract 8 from both sides to get:
3x = 6
Divide both sides by 3 to get:
x = 2
================================
Cheers,
Stan H.